# Is Noise Conditioning Necessary for Denoising Generative Models?

### v1

@@ Appendix E Additional Samples
Beyond the comparison shown in Fig. 5 for noise-conditional and noise-unconditional models, we also provide additional samples for other models, on other datasets, or in class-conditional settings. We use the same configuration as in Table 2. Figs. 9 and 10 show the samples of ICM and ECM on CIFAR-10 with both 1 and 2 inference steps. Fig. 11 show the samples of FM on ImageNet 32$\times$32 with both Euler and EDM-Heun sampler. Fig. 12 shows the samples of FM and EDM on CIFAR-10 in a class-conditional setting.
@@
Figure 11: Samples generated by FM on ImageNet 32$\times$32 with Euler and EDM-Heun sampler. From left to right: Euler w/ $t$, Euler w/o $t$, Heun w/ $t$, Heun w/o $t$. All corresponding samples use the same noise.

### v2

@@ Appendix E Additional Samples
Beyond the comparison shown in Fig. 5 for noise-conditional and noise-unconditional models, we also provide additional samples for other models, on other datasets, or in class-conditional settings. We use the same configuration as in Table 2. Figs. 10 and 11 show the samples of ICM and ECM on CIFAR-10 with both 1 and 2 inference steps. Fig. 12 show the samples of FM on ImageNet $\times32$ with both Euler and EDM-Heun sampler. Fig. 13 shows the samples of FM and EDM on CIFAR-10 in a class-conditional setting.
@@
Figure 12: Samples generated by FM on ImageNet $\times32$ with Euler and EDM-Heun sampler. From left to right: Euler w/ $t$, Euler w/o $t$, Heun w/ $t$, Heun w/o $t$. All corresponding samples use the same noise.